OpenStudy (help_people):

The employees of two companies were asked how many times they checked their email during weekends. The table below shows the numbers for each company: Company A 2 4 8 8 2 6 4 8 3 Company B 2 6 3 5 2 4 6 3 2 Based on the table, which of the following is true? The interquartile range for Company A employees is 2 more than the interquartile range for Company B employees. The interquartile range for Company A employees is 0.5 more than the interquartile range for Company B employees. The interquartile range for Company A employees is 1 more than the interquartile range for Company

OpenStudy (jack1):

hey so first: do you understand what an interquartile range is? if so please define it below

OpenStudy (jack1):

better yet: what is the interqurtile range for company A and for company B

OpenStudy (help_people):

hi

OpenStudy (help_people):

and i will try is it when ur dividing the data set in qurtiles examples would be q1 q2?

OpenStudy (jack1):

yep, divide the data set for each company into 3 parts first then find ur median then IQR = Q3 - Q1

OpenStudy (help_people):

imnot sure hiw to do that! :D

OpenStudy (jack1):

ok company A ill do, company B u try from my example, yeah?

OpenStudy (help_people):

ok

OpenStudy (jack1):

Company A 2 4 8 8 2 6 4 8 3 so 9 numbers, so lest group them by 3's lowest 3 = 2,2,3 => Q1 = 2.5 middle 3 = 4,4,6 => Q2 = 4 upper 3 = 8,8,8 => Q3 = 8 median number = 4 there are 4 numbers on each side of the median number so IQR = Q3 - Q1 = 8 - 2.5 = 5.5 ... i think? my stats are rusty though

OpenStudy (help_people):

ok

OpenStudy (texaschic101):

put the numbers in order... 2,2,3,4,4,6,8,8,8 Q2 (median--middle number) = 4 Q1 = (2 + 3)/2 = 6/2 = 3 Q3 = (8 + 8)/2 = 16/2 = 8 IQR = Q3 - Q1 IQR = 8 - 3 = 5

OpenStudy (texaschic101):

I am getting something different for the first one

OpenStudy (jack1):

3+2 = 5

OpenStudy (jack1):

not 6

OpenStudy (texaschic101):

oops..your right Q1 = (3 + 2)/2 = 5/2 = 2.5....your right so IQR is 8 - 2.5 = 5.5

OpenStudy (texaschic101):

good catch Jack...I am feeling real stupid right now...lol

OpenStudy (jack1):

hehe, s'cool, we all have brain hiccups sometimes ;)

OpenStudy (jack1):

ok @help_people now u try with company B example pic attached to assist if my example didnt help that much?

OpenStudy (texaschic101):

you might have to explain what to do if there is 2 middle numbers

OpenStudy (texaschic101):

because of the quartiles

OpenStudy (jack1):

oh yeah, soz um, if u find the median number, and there's an odd number either side, the quartiles are the middle of those odd numbers eg 7 numbers 1,2,3,4,5,6,7 median = 4 3 numbers either side 1,2,3 4 5,6,7 middle number if each side is the quartile 1,2,3 ... Q1 = 2 5,6,7 ... Q3 = 6

OpenStudy (help_people):

Q1=2 Q2=3 Q3=6

OpenStudy (help_people):

for company B

OpenStudy (jack1):

now if u have an even number either side 1,2,3,4,5,6,7,9,15 median = 5 4 numbers either side 1,2,3,4 5 6,7,9,15 so to find quartile, you need to find the middle number of each side 1,2,3,4 middle numbers are 2 and 3 middle of this is 2.5 =Q1 6,7,9,15 .... middle numbers are 7 and 9 middle of this = 8 = Q3

OpenStudy (jack1):

Company B 2 6 3 5 2 4 6 3 2 2,2,2,3,3,4,5,6,6 Q1 = 2 Q2 = 3 Q3 = 5+6/2 = 5.5

OpenStudy (texaschic101):

^^ I agree 2,2,2,3,3,4,5,6,6 Q2 = 3 Q1 = (2+2)/2 = 4/2 = 2 Q3 = (5 + 6) / 2 = 11/2 = 5.5 IQR = 5.5 - 2 = ?

OpenStudy (help_people):

oh srry 3.5

OpenStudy (help_people):

my internet isreally slow where im ay so bear with me

OpenStudy (texaschic101):

so A = 5.5 and B = 3.5.....so what do you think the answer is now

OpenStudy (help_people):

a is the answer

OpenStudy (texaschic101):

yep ^^ and be sure and give Jack a medal :)

OpenStudy (jack1):

sorry, loo break, yeps A is the best ;)

OpenStudy (jack1):

cheers for the assist @texaschic101 i wasn't 100% that i remembered IQR's right till u came along